OpenStudy (anonymous):

Which is an equation in point-slope form of the line that passes through the points (7,5) and (-4,-1)?

undeadknight26 (undeadknight26):

to find the oint slope form we must first find the slope. Do you know how to find the slope?

OpenStudy (anonymous):

no

undeadknight26 (undeadknight26):

y2 - y1/x2 - x1 -1 - 5/-4 - 7 = m or slope.

OpenStudy (anonymous):

-6/-11

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

anyone there

OpenStudy (anonymous):

Y-Y1=(Y2-Y1)(X-X1) / (X2-X1)

OpenStudy (anonymous):

where (X1,Y1) = (7,5) and (X2,Y2) = (-4,-1).Substitute in the given equation,the u get the equation of line.

OpenStudy (anonymous):

y=7-5(x-(-4))-(-1)